# Unraveling the Enigma of Double Descent: An In-depth Analysis through the Lens of Learned Feature Space

### v1

@@ A.2 Experiment Setups
  • Optimization: For all our experiments, we employ the Stochastic Gradient Descent (SGD) optimizer with zero momentum and a batch size of 128. The learning rate starts with an initial value of 0.05 and updates every 50 epochs. Below, we provide a precise description of our different learning rate schedule on SimpleFC and CNNs/ResNet18s:
@@ A.3.1 Model Performance in the Presence of High Label Noise
We provided here two additional experiment results when a high label noise ratio $p=40\%$ is introduced to the training dataset of MNIST and CIFAR-10. It is evident that noise has a substantial impact on the model’s classification accuracy, leading to a reduction. However, the correlation between ${\rm Kp}$ and generalization performance remains consistent.
@@
Figure 8: The phenomenon of double descent when a high label noise of $p=40\%$ is introduced on the training dataset.
A.3.2 t-SNE visualization of Learned Representations of trained ResNet18s
We provided here the t-SNE visualization of Learned Representations of trained ResNet18s mapping from a 1000 random samples from the training dataset of CIFAR-10.
Figure 9: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained ResNet18s of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models are able to segregate noise within the learned feature space.

### v2

@@ A.2 Experiment Setups
  • Optimization: For all our experiments, we employ the Stochastic Gradient Descent (SGD) optimizer with zero momentum and a batch size of 128. The learning rate starts with an initial value of 0.05 and updates every 50 epochs. Below, we provide a precise description of our different learning rate schedules on SimpleFC and CNNs/ResNet18s:
@@ A.3.1 Model Performance in the Presence of High Label Noise
We provided here two additional experiment results when a high label noise ratio $p=40\%$ is introduced to the training dataset of MNIST and CIFAR-10. It is evident that noise has a substantial impact on the model’s classification accuracy, leading to a reduction. However, the correlation between $P$ and generalization performance remains consistency.
@@
Figure 6: The phenomenon of double descent and the prediction accuracy $P$ when a high label noise of $p=40\%$ is introduced on the training dataset.
A.3.2 t-SNE visualization of Learned Representations
In addition to evaluating the value of $P$, we also utilize T-Distributed Stochastic Neighbor Embedding (t-SNE) to provide a visual representation of the learned features in the feature space. While this visualization may not accurately represent the intricate inherent structure of high-dimensional features, it aims to gain insights into the clustering and distribution of data points, thereby enhancing our understanding of the model’s internal representations.
The t-SNE visualization in Figure.7 offers evidence from a more intuitive standpoint: as layer width increases, the overlap between different classes diminishes, and noisy data points are more accurately assigned to their correct classes. The choice of layer width $k=[10,20,100]$ represents three different circumstances. When $k=10$, the bias-variance trade-off is reached at the bottom of the U-shaped segment of the test error curve. The training dataset is not fully interpolated and the learned representations are less organized. When $k=20$, the model is about to interpolate the whole training set, which includes the noisy labeled data. From the figure, it’s evident that at this stage, representations of some data points with noisy labels appear within classes they don’t belong to, which are likely to introduce interference and lead to misleading effects. In the case of an over-parameterized model, such as when $k=100$, this phenomenon significantly diminishes, and the majority of noisy labeled data points are encoded alongside their true classmates.
Figure 7: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained SimpleFC of 1000 randomly selected data points from the training set of MNIST, at different layer widths of $k=[10,20,100]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models can segregate noise within the learned feature space.
When we examine the t-SNE visualization of learned representations in CNN trained on the CIFAR-10 dataset, despite the increased complexity of 2D distributions due to the higher dimensionality of CIFAR-10 images, we can still observe distinct class distributions and improved identification of noisy data points within these distributions as the model width expands. The t-SNE visualization Figure.9 of ResNets conveys the same statement.
Figure 8: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained CNNs of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, indicating that over-parameterized models can segregate noise within the learned feature space.
Figure 9: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained ResNet18s of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models can segregate noise within the learned feature space.